OpenStudy (anonymous):

find the slope of the line passing through the points (1980, 16000) (2000, 42000)

OpenStudy (anonymous):

these are big numbers, but just remember, slope is the difference in the y values divided by the difference in the x values. slope = (y2 - y1) / (x2 - x1) between points (x1,y1) and (x2,y2)

OpenStudy (anonymous):

I thought that was the way to come up with the solution; thanks!

OpenStudy (anonymous):

no problem :) so it's (42000 - 16000) / (2000 - 1980) = 26000 / 20 = 1300 Is that what you get?

OpenStudy (anonymous):

Yes, it sure is :)

OpenStudy (anonymous):

Great :)